A bicyle shop has bicycles with 2 wheels and tricycles with 3 wheels. They have 15 different cycles for sale. The 15 cycles have 34 wheels. How many bicycles and tricycles does the shop have for sale?
Found 2 solutions by solver91311, Greenfinch:
Answer by solver91311(24713)



Let represent the number of 2 wheeled cycles.

Let represent the number of 3 wheeled cycles.

Since the shop had 15 cycles and they are all either 2 wheeled or 3 wheeled, then:



The number of wheels on two-wheeled cycles must be . The number of wheels on three-wheeled cycles must be . And the sum of the wheels is 34, so:



Just solve the system of equations and you have your answer.

John


Answer by Greenfinch(383)

Let the number of bicycles be B
Then the number of tricycles is 15 - B
No of wheels is thus 2B + 3(15 - B) = 34
2B + 45 -3B = 34
-B = - 11
So there are 11 bicycles and therefore 4 tricycles